Let V be a vector space of 2×2 matrices over R. Show that the set S defined by S={(a,b)(c,d)belongs to V :a+b=0} is a subspace of R

1
Expert's answer
2022-02-18T13:04:11-0500
  • First of all, let us remark that the zero matrix satisfies the property "a+b=0" so "0\in S"

Let "M, N" be matrices in "S" and "\lambda" "\in \mathbb{R}" a scalar. We will denote "a_M" the top-left coefficient of a matrix "M" and "b_M" the top-right one (as in the question).

  • By definition of a sum of two matrices, we have "a_{M+N}=a_M+a_N", same for "b_{M+N}", so we have "a_{M+N}+b_{M+N}=(a_M+b_M)+(a_N+b_N)=0" and thus "M+N\in S"
  • By definition of a scalar multiplication we have "a_{\lambda M} = \lambda a_M", same for "b_{\lambda M}" and thus we have "a_{\lambda M}+b_{\lambda M}= \lambda(a_M+b_M)=0" and thus "\lambda M\in S".

Therefore, "S" is a vector subspace of "V".